• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name G. M. Urciuoli

    Log entry time 08:11:43 on November 1,2001

    Entry number 75540

    keyword=11/1/2001 owl shift summary


    Day: 11/01/2001 Time: 00:00
    Owl Shift Starts

    Shift Crew: G. M. Urciuoli
    Issam Qattan

    Run N. of Events Sum_Q 1p1/2 1p3/2
    (uA-hours)
    ***** ********* ********** ******* *********
    1405 1.5 3384 440 2022

    1406 1.5 3467 422 2069

    1407 1.5 3553 348 1889

    1408 1.5 3642 404 1638

    1409 1.5 3732 291 1514

    1410 1.5 3817 197 1126

    1411 0.2 3827 22 80

    (1) During the 1410 run, the left arm Q1 current did not stay in lock for a while. The problem was fixed by lowering the current manually to 3100.00 uA.
    (2) During the 1411 run, beam was turned off at 5:35 AM . We stopped running.
    (3) At 6:00 AM, the MCC called to put the fast feedback off (Masked). The beam set energy was = 4620.07 MeV, and Tief Energy was = 4617.93 MeV.
    (4) At 7:00 AM, the target was removed.
    (5) At 7:30 AM, the target was back on.
    (6) At 7:35 AM, the target was removed again.